# Dyson Equation for Correlated Linearizations and Test Error of Random Features Regression

### v2

@@ 3.3 Gaussian Equivalence
As a consequence of Theorem 2.4, we discuss in Section 2.4.5 that random features ridge regression follows a Gaussian equivalence principle. This means that in terms of training and empirical test error, it behaves equivalently to a surrogate linear Gaussian model with matching covariance. This phenomenon was initially established for random features ridge regression under the linear regime in the sense of test error [51]. Subsequent work proved its extension to broader loss functions and regularization, initially via non-rigorous replica methods [29] and later through rigorous analysis [31, 40]. Gaussian equivalence has also been demonstrated for deep random features [56] and for random features ridge regression beyond the linear scaling regime [41].
Our work extends this literature by establishing Gaussian equivalence in terms of the empirical test error of random features ridge regression. This contributes to the generalization of Gaussian equivalence under broader distributional assumptions, aligning with the direction explored by [57].
@@ 3.4 Training and Test Error of Random Features
The random features model provides valuable insights into the behavior of more complex machine learning models and serves as a useful benchmark. Previous studies have extensively examined the test error of random features ridge regression in the proportional isotropic regime both using non-rigorous replica methods [29] and rigorous analyses [51, 1, 3]. This line of research characterizes the training and test error of random features ridge regression in order to offer insights into the impacts of various model choices, such as overparameterization.
Research has expanded to address anisotropic data [36, 52, 50, 57] and covariate shift scenarios [59]. Beyond the linear scaling regime, studies revealed significant transitions in the degree of label learning as a function of the polynomial scaling exponent [41, 50]. Additionally, some studies investigated training, test, and cross-validation errors in the highly overparameterized regime with nearly orthogonal data assumptions [62]. Another line of research investigates the test error of random features beyond ridge regression. Studies have explored generic convex losses [31] and alternative penalty terms [47, 19]. A recent trend focuses on deep random features, a multi-layer generalization of random features. Empirical findings indicate that trained neural network outputs can be modeled by a deep random features model, with each layer’s covariance corresponding to that of the neural network [32]. Research has delved into deep structured linear networks [64] and deep non-linear networks [19, 56, 57].
Our study diverges from these works by focusing on the empirical test error of random features ridge regression, without assuming specific data models or distributions beyond some boundedness conditions. Our main result regarding the test error of random features ridge regression is most similar to the work of [46], who established an asymptotically exact expression for the training error of random features ridge regression. As previously mentioned, the authors conjectured that Theorem 2.4 holds without the additional conditions imposing bounded expectations for the norm of the random features matrices. [45] resolves this conjecture in the special case of random Fourier features. Both [45, 46] employ leave-one-out techniques and concentration of measure arguments in their approaches, as do we in part of the universality argument.
@@ 4 Properties of the Dyson Equation for Linearizations
assuming its well-definedness, which we establish through lemmas 4.1, 4.3 and 4.4. With this definition, we can reexpress (2) as $M=\mathcal{F}(M)$. Whenever convenient, we will fix a spectral parameter $z\in\mathbb{H}$ and operate with $\mathcal{F}$ over $\mathcal{A}$. Similarly, the formulation of (5) becomes $M^{(\tau)}=\mathcal{F}^{(\tau)}(M^{(\tau)})$, where
@@ 4.1 General Properties
Let $M\in\mathcal{A}$ be arbitrary. By definition of $\mathcal{A}$, $\Im[M]\succeq 0$. Since $\mathcal{S}$ is symmetric, linear and positivity-preserving, $\Im[\mathcal{S}(M)]=\mathcal{S}(\Im[M])\succeq 0$. If $\tau>0$, it follows directly that $\Im[\mathbb{E}L-\mathcal{S}(M)-z\Lambda-i\tau I_{\ell}]\preceq-\tau$, which implies that $\mathbb{E}L-\mathcal{S}(M)-z\Lambda-i\tau I_{\ell}$ is non-singular by Lemma A.6.
@@
Again, since the real and imaginary parts of a matrix are Hermitian, the quadratic forms are real and $v_{2}^{\ast}\Im[\mathbb{E}Q-\mathcal{S}_{2,2}(M)]v_{2}=v_{2}^{\ast}\Re[\mathbb{E}Q-\mathcal{S}_{2,2}(M)]v_{2}=0$. In particular, since $M\in\mathcal{A}$, $\Im[M_{1,1}]\succ 0$ and $0=v_{2}^{\ast}\Im[\mathbb{E}Q-\mathcal{S}_{2,2}(M)]v_{2}\leq-v_{2}^{\ast}\mathcal{S}_{2,2}(\Im[M])v_{2}\leq 0$. By definition of $\mathcal{S}_{2,2}$, $v_{2}^{\ast}\mathcal{S}_{2,2}(\Im[M])v_{2}=\mathbb{E}v^{\ast}_{2}(B-\mathbb{E}B)\Im[M_{1,1}](B-\mathbb{E}B)^{T}v_{2}$. Consequently, as $\Im[M_{1,1}]$ is positive definite, it must be the case that $(B-\mathbb{E}B)^{T}v_{2}=0$ almost surely. Going back to the equation $(\mathbb{E}Q-\mathcal{S}_{2,2}(M))v_{2}=0$, we obtain $\mathbb{E}Qv_{2}=0$. However, $\mathbb{E}Q$ is non-singular, which implies that $v_{2}=0$. This is a contradiction, the kernel of $\mathbb{E}L-\mathcal{S}(M)-z\Lambda$ is trivial, and the matrix is non-singular. ∎
For every $M\in\mathcal{A}$, $\tau\in\mathbb{R}_{\geq 0}$ and $z\in\mathbb{H}$, the upper-left $n\times n$ block of $\mathbb{E}L-\mathcal{S}(M)-z\Lambda-i\tau I_{\ell}$ has negative definite imaginary part. Hence, by Lemma A.6, the upper-left $n\times n$ block is non-singular. By Lemma A.10, this implies that the full matrix $\mathbb{E}L-\mathcal{S}(M)-z\Lambda-i\tau I_{\ell}$ is non-singular if and only if the associated Schur complement is non-singular. Since we established non-singularity of the full matrix, this implies that the lower-right block of $(\mathbb{E}L-\mathcal{S}(M)-z\Lambda-i\tau I_{\ell})^{-1}$ is non-singular. Similarly, the previous lemma establishes that the lower-right $d\times d$ block of $\mathbb{E}L-\mathcal{S}(M)-z\Lambda-i\tau I_{\ell}$ is non-singular. Additionally, the associated Schur complement has a negative definite imaginary part, which, by Lemma A.6, implies that the Schur complement is non-singular. Consequently, we obtain the following corollary.
@@
Lemma 4.1 is a first step towards considering the DEL (2) as a fixed point equation $\mathcal{F}(M)=M$, along with its regularized counterpart. A second step in this direction is showing that $\mathcal{F}$ and $\mathcal{F}^{(\tau)}$ both map their respective domains to themselves. We adapt the argument from [37].
@@
Combining lemmas 4.1 and 4.3, it only remains to show that the DEL map $\mathcal{F}$ and the RDEL map $\mathcal{F}^{(\tau)}$ preserve holomorphicity to establish that they map their respective domains to themselves.
@@
The insights garnered from lemmas 4.4, 4.3 and 4.1 leads us to conceptualize the DEL and RDEL as fixed point equations. This characterization is instrumental, as we rely on it to assert the existence of a unique solution for (5) using the contractive property of the RDEL map $\mathcal{F}^{(\tau)}$ with respect to the CRF-pseudometric. Furthermore, this perspective will be crucial to establish the stability of the DEL.
@@ 4.2 Large Spectral Parameter
and denote $M_{\star}=M_{\star}^{(0)}$, $M_{\infty}=M_{\infty}^{(0)}$. Indeed, by Lemma A.7, $\Im[M^{(\tau)}_{\ast}]\succeq 0$. We demonstrate in Lemma 4.5 that $M^{(\tau)}_{\star}$ corresponds precisely to the limit of $M^{(\tau)}(z)$ as $\Im[z]$ diverges to infinity.
@@
Fix $\tau\in\mathbb{R}_{\geq 0}$ and assume that $M\in\mathcal{M}$ such that, for all $z\in\mathbb{H}$, $M(z)$ solves the RDEL (5). Then, $\|M(z)-M_{\infty}\|\to 0$ as $\Im[z]\to\infty$.
@@
We proceed block-wise. By Lemma 4.3, $\|M_{1,1}(z)\|\leq(\Im[z])^{-1}$. Consequently, $\|M_{1,1}(z)\|\to 0$ as $\Im[z]\to\infty$. Furthermore, it follows from Lemma A.7 and the assumed properties of the superoperator that $\Im[(\mathcal{T}^{(\tau)}_{1,1}(M))^{-1}]\succeq\Im[z](\mathcal{T}^{(\tau)}_{1,1}(M))^{-1}(\mathcal{T}^{(\tau)}_{1,1}(M))^{-\ast}$. Here, we use the notation $\mathcal{T}^{(\tau)}(M)=\mathbb{E}L-\mathcal{S}(M)-z\Lambda-i\tau I_{\ell}$ introduced in Equation 14. Taking the norm of both sides and rearranging, we obtain $\|(\mathcal{T}^{(\tau)}_{1,1}(M))^{-1}\|\leq(\Im[z])^{-1}$. Hence, $\|(\mathcal{T}^{(\tau)}_{1,1}(M))^{-1}\|\to 0$ as $\Im[z]\to\infty$. Furthermore, using the flatness of the superoperator, we have $\|\mathcal{S}_{1,2}(M)\|\vee\|\mathcal{S}_{2,1}(M)\|\vee\|\mathcal{S}_{2,2}(M)\|\to 0$ as $\Im[z]\to\infty$. This implies that $\mathcal{T}_{1,2}(M)\to\mathbb{E}B^{T}$, $\mathcal{T}_{2,1}(M)\to\mathbb{E}B$ and $\mathcal{T}_{2,2}(M)\to\mathbb{E}Q-i\tau I_{d}$ as $\Im[z]\to\infty$. Since $\mathbb{E}Q-i\tau I_{d}$ is non-singular and the taking a matrix inverse is a continuous operation, $(\mathcal{T}_{2,2}(M))^{-1}\to M^{(\tau)}_{\ast}$ as $\Im[z]\to\infty$. Finally, using Equation 14, we conclude that $M(z)\to M_{\infty}$ as $\Im[z]\to\infty$. ∎
@@
Similarly to Lemma 4.5, it follows from lemmas A.10 and A.6 that $\|(L-z\Lambda-i\tau I_{\ell})^{-1}-\mathrm{diag}\{0_{n\times n},(Q-i\tau I_{d})^{-1}\}\|$ as $\Im[z]\to\infty$.
@@
Fix $\tau\in\mathbb{R}_{\geq 0}$ and assume that $M\in\mathcal{M}$ such that, for all $z\in\mathbb{H}$, $M(z)$ solves the RDEL (5). Then, there exists some constant $c\in\mathbb{R}_{>0}$ such that $\|M(z)-M^{(\tau)}_{\infty}\|\leq c(|z|-\kappa)^{-1}$ for all $z\in\{z\in\mathbb{H}\,:\,|z|>\kappa+c\kappa^{-1}\}$ with $\kappa:=2\|(\mathbb{E}Q-i\tau I_{d})^{-1}\|(\|\mathbb{E}B\|+(2\|(\mathbb{E}Q-i\tau I_{d})^{-1}\|)^{-1})^{2}+\|\mathbb{E}A\|+(2\|(\mathbb{E}Q-i\tau I_{d})^{-1}\|)^{-1}+s\|(\mathbb{E}Q-i\tau I_{d})^{-1}\|$.
@@
Fix $z\in\mathbb{H}$ with $|z|>\kappa$ and let $M\equiv M(z)$. For notational convenience, we denote $a=\|\mathbb{E}A\|$, $b=\|\mathbb{E}B\|$ and $m_{\star}=\|M_{\star}^{(\tau)}\|$. We will show that there exists $c\in\mathbb{R}_{>0}$ such that $\|M(z)-M_{\infty}^{(\tau)}\|\notin(c(|z|-\kappa)^{-1},\kappa]$ for every $z\in\mathbb{H}$ with $|z|>\kappa+c\kappa^{-1}$. By Lemma 4.5, $\|M(z)-M_{\infty}^{(\tau)}\|$ is in a neighborhood of the origin for every $z\in\mathbb{H}$ with $\Im[z]$ large enough. Since $z\mapsto\|M(z)-M_{\infty}^{(\tau)}\|$ is a continuous function on $\{z\in\mathbb{H}\,:\,|z|>\kappa+c\kappa^{-1}\}$, this will imply that $\|M(z)-M_{\infty}^{(\tau)}\|\in[0,c(|z|-\kappa)^{-1}]$ for every $z\in\{z\in\mathbb{H}\,:\,|z|>\kappa+c\kappa^{-1}\}$.
Suppose that $\|M-M^{(\tau)}_{\infty}\|\leq(2sm_{\star})^{-1}$ such that $\|M\|\leq\|M-M^{(\tau)}_{\infty}\|+\|M^{(\tau)}_{\infty}\|\leq(2sm_{\star})^{-1}+m_{\star}$. We consider the blocks separately using (14). By definition of $\mathcal{S}_{2,2}$, $\|\mathcal{S}_{2,2}(M)\|\leq s\|M_{1,1}\|\leq(2m_{\star})^{-1}$. It follows from Lemma A.3 that
$\|(\mathcal{T}_{2,2}(M))^{-1}\|=\|(\mathbb{E}Q-i\tau I_{d})^{-1}\left(\mathcal{S}_{2,2}(M)(\mathbb{E}Q-i\tau I_{d})^{-1}-I_{d}\right)^{-1}\|\leq 2m_{\star}.$
@@
$\|\mathbb{E}A-\mathcal{S}_{1,1}(M)-\mathcal{T}_{1,2}(M)(\mathcal{T}_{2,2}(M))^{-1}\mathcal{T}_{2,1}(M)\|\leq a+(2m_{\star})^{-1}+2m_{\star}(b+(2m_{\star})^{-1})^{2}.$
For $z\in\mathbb{H}$ with $|z|>\kappa\geq a+(2m_{\star})^{-1}+2m_{\star}(b+(2m_{\star})^{-1})^{2}$, it follows from Lemmas A.3 and 14a that $\|M_{1,1}\|\leq(|z|-\kappa)^{-1}$. We now turn our attention to $M_{2,2}$. Using Lemmas A.2 and 14b,
$M_{2,2}-M_{\star}^{(\tau)}=M_{2,2}\left(\mathcal{S}_{2,2}(M)+\mathcal{T}_{2,1}(M)(\mathcal{T}_{1,1}(M))^{-1}\mathcal{T}_{1,2}(M)\right)^{-1}M^{(\tau)}_{\ast}.$
By Lemma A.3, $\|(\mathcal{T}_{1,1}(M))^{-1}\|\leq(|z|-a-(2m_{\star})^{-1}-sm_{\star})^{-1}$. Hence,
$\|M_{2,2}-M_{\star}^{(\tau)}\|\leq sm_{\star}((2sm_{\star})^{-1}+m_{\star})\|M_{1,1}\|\\ +m_{\star}((2sm_{\star})^{-1}+m_{\star})(b+(2m_{\star})^{-1})^{2}(|z|-\kappa)^{-1}.$
Plugging the bound for $\|M_{1,1}\|$ derived above and simplifying,
$\|M_{2,2}-M_{\star}^{(\tau)}\|\leq m_{\star}(s+(b+(2m_{\star})^{-1})^{2})((2sm_{\star})^{-1}+m_{\star})(|z|-\kappa)^{-1}.$
It only remains to treat $\|M_{1,2}\|$ and $\|M_{2,1}\|$, which we directly bound by
$\max\{\|M_{1,2}\|,\|M_{2,1}\|\}\leq 2m_{\star}(b+(2m_{\star})^{-1})\|M_{1,1}\|\leq 2m_{\star}(b+(2m_{\star})^{-1})(|z|-\kappa)^{-1}$
using (14c) and (14d). To summarize, we showed that for every $z\in\mathbb{H}$ with $|z|>a+(2m_{\star})^{-1}+2m_{\star}(b+(2m_{\star})^{-1})^{2}+sm_{\star}$, $\|M(z)-M^{(\tau)}_{\infty}\|\leq(2sm_{\star})^{-1}$ implies that
$\|M(z)-M^{(\tau)}_{\infty}\|\leq\|M_{1,1}(z)\|+\|M_{1,2}\|+\|M_{2,1}\|+\|M_{2,2}-M^{(\tau)}_{\star}\|\leq c(|z|-\kappa)^{-1}$

### v4

@@ 3.3 Gaussian Equivalence
As a consequence of Theorem 2.3, we discuss in Section 2.4.5 that random features ridge regression follows a Gaussian equivalence principle. This means that in terms of training and empirical test error, it behaves equivalently to a surrogate linear Gaussian model with matching covariance. This phenomenon was initially established for random features ridge regression under the linear regime in the sense of test error [montanari_RF]. Subsequent work proved its extension to broader loss functions and regularization, initially via non-rigorous replica methods [pmlr-v119-gerace20a] and later through rigorous analysis [goldt2022gaussian, hong_lu_GET]. Gaussian equivalence has also been demonstrated for deep random features [schroder2023deterministic] and for random features ridge regression beyond the linear scaling regime [hu2024asymptotics].
Our work extends this literature by establishing Gaussian equivalence in terms of the empirical test error of random features ridge regression. This contributes to the generalization of Gaussian equivalence under broader distributional assumptions, aligning with the direction explored by [schröder2024asymptotics].
@@ 3.4 Training and Test Error of Random Features
The random features model provides valuable insights into the behavior of more complex machine learning models and serves as a useful benchmark. Previous studies have extensively examined the test error of random features ridge regression in the proportional isotropic regime both using non-rigorous replica methods [pmlr-v119-gerace20a] and rigorous analyses [montanari_RF, Adlam2019ARM, adlam_doubledescent]. This line of research characterizes the training and test error of random features ridge regression in order to offer insights into the impacts of various model choices, such as overparameterization.
Research has expanded to address anisotropic data [hastie_22, mel2022anisotropic, MEI20223, schröder2024asymptotics] and covariate shift scenarios [tripuraneni2021overparameterization]. Beyond the linear scaling regime, studies revealed significant transitions in the degree of label learning as a function of the polynomial scaling exponent [hu2024asymptotics, MEI20223]. Additionally, some studies investigated training, test, and cross-validation errors in the highly overparameterized regime with nearly orthogonal data assumptions [wang2023overparameterized]. Another line of research investigates the test error of random features beyond ridge regression. Studies have explored generic convex losses [goldt2022gaussian] and alternative penalty terms [Loureiro_2022, bosch2023precise]. A recent trend focuses on deep random features, a multi-layer generalization of random features. Empirical findings indicate that trained neural network outputs can be modeled by a deep random features model, with each layer’s covariance corresponding to that of the neural network [guth2023rainbow]. Research has delved into deep structured linear networks [zavatone_deep_linear] and deep non-linear networks [bosch2023precise, schroder2023deterministic, schröder2024asymptotics].
Our study diverges from these works by focusing on the empirical test error of random features ridge regression, without assuming specific data models or distributions beyond some boundedness conditions. Our main result regarding the test error of random features ridge regression is most similar to the work of [louart_random-matrix-approach], who established an asymptotically exact expression for the training error of random features ridge regression. As previously mentioned, the authors conjectured that Theorem 2.3 holds without the additional conditions imposing bounded expectations for the norm of the random features matrices. [Liao_2021] resolves this conjecture in the special case of random Fourier features. Both [Liao_2021, louart_random-matrix-approach] employ leave-one-out techniques and concentration of measure arguments in their approaches, as do we in part of the universality argument.
@@ 4 Properties of the Dyson Equation for Linearizations
assuming its well-definedness, which we establish through lemmas 4.1, 4.2 and 4.3. With this definition, we can reexpress (2) as $M=\mathcal{F}(M)$. Whenever convenient, we will fix a spectral parameter $z\in\mathbb{H}$ and operate with $\mathcal{F}$ over $\mathcal{A}$. Similarly, the formulation of (5) becomes $M^{(\tau)}=\mathcal{F}^{(\tau)}(M^{(\tau)})$, where
@@ 4.1 General Properties
Let $M\in\mathcal{A}$ be arbitrary. By definition of $\mathcal{A}$, $\Im[M]\succeq 0$. Since positivity-preserving, we may use Lemma A.11 to get $\Im[\mathcal{S}(M)]=\mathcal{S}(\Im[M])\succeq 0$. If $\tau>0$, it follows directly that $\Im[\mathbb{E}L-\mathcal{S}(M)-z\Lambda-i\tau I_{\ell}]\preceq-\tau$, which implies that $\mathbb{E}L-\mathcal{S}(M)-z\Lambda-i\tau I_{\ell}$ is non-singular by Lemma A.5.
@@
Again, since the real and imaginary parts of a matrix are Hermitian, the quadratic forms are real and $v_{2}^{\ast}\Im[\mathbb{E}Q-\mathcal{S}_{2,2}(M)]v_{2}=v_{2}^{\ast}\Re[\mathbb{E}Q-\mathcal{S}_{2,2}(M)]v_{2}=0$. In particular, since $M\in\mathcal{A}$, $\Im[M_{1,1}]\succ 0$ and $0=v_{2}^{\ast}\Im[\mathbb{E}Q-\mathcal{S}_{2,2}(M)]v_{2}\leq-v_{2}^{\ast}\mathcal{S}_{2,2}(\Im[M])v_{2}\leq 0$, where we have once again used Lemma A.11. By definition of $\mathcal{S}_{2,2}$, $v_{2}^{\ast}\mathcal{S}_{2,2}(\Im[M])v_{2}=\mathbb{E}v^{\ast}_{2}(B-\mathbb{E}B)\Im[M_{1,1}](B-\mathbb{E}B)^{\top}v_{2}$. Consequently, as $\Im[M_{1,1}]$ is positive definite, it must be the case that $(B-\mathbb{E}B)^{\top}v_{2}=0$ almost surely. Going back to the equation $(\mathbb{E}Q-\mathcal{S}_{2,2}(M))v_{2}=0$, we obtain $\mathbb{E}Qv_{2}=0$. However, $\mathbb{E}Q$ is non-singular, which implies that $v_{2}=0$. This is a contradiction, the kernel of $\mathbb{E}L-\mathcal{S}(M)-z\Lambda$ is trivial, and the matrix is non-singular. ∎
For every $M\in\mathcal{A}$, $\tau\in\mathbb{R}_{\geq 0}$ and $z\in\mathbb{H}$, the upper-left $n\times n$ block of $\mathbb{E}L-\mathcal{S}(M)-z\Lambda-i\tau I_{\ell}$ has negative definite imaginary part. Hence, by Lemma A.5, the upper-left $n\times n$ block is non-singular. By Lemma A.9, this implies that the full matrix $\mathbb{E}L-\mathcal{S}(M)-z\Lambda-i\tau I_{\ell}$ is non-singular if and only if the associated Schur complement is non-singular. Since we established non-singularity of the full matrix, this implies that the lower-right block of $(\mathbb{E}L-\mathcal{S}(M)-z\Lambda-i\tau I_{\ell})^{-1}$ is non-singular. Similarly, the previous lemma establishes that the lower-right $d\times d$ block of $\mathbb{E}L-\mathcal{S}(M)-z\Lambda-i\tau I_{\ell}$ is non-singular. Additionally, the associated Schur complement (see Lemma A.9) has a negative definite imaginary part, which, by Lemma A.5, implies that the Schur complement is non-singular. Consequently, we obtain the following corollary.
@@
Lemma 4.1 is a first step towards considering the DEL (2) as a fixed point equation $\mathcal{F}(M)=M$, along with its regularized counterpart. A second step in this direction is showing that $\mathcal{F}$ and $\mathcal{F}^{(\tau)}$ both map their respective domains to themselves. We adapt the argument from [helton2007operatorvalued].
@@
Combining lemmas 4.1 and 4.2, it only remains to show that the DEL map $\mathcal{F}$ and the RDEL map $\mathcal{F}^{(\tau)}$ preserve holomorphicity to establish that they map their respective domains to themselves.
@@
The insights garnered from lemmas 4.3, 4.2 and 4.1 leads us to conceptualize the DEL and RDEL as fixed point equations. This characterization is instrumental, as we rely on it to assert the existence of a unique solution for (5) using the contractive property of the RDEL map $\mathcal{F}^{(\tau)}$ with respect to the CRF-pseudometric. Furthermore, this perspective will be crucial to establish the stability of the DEL.
@@ 4.2 Large Spectral Parameter
and denote $_{\star}=M_{\star}^{(0)}, M=M$$_{\infty}=M_{\infty}^{(0)}$. Indeed, by Lemma A.6, $\Im[M^{(\tau)}_{\ast}]\succeq 0$. We demonstrate in Lemma 4.4 that $M^{(\tau)}_{\star}$ corresponds precisely to the limit of $M^{(\tau)}(z)$ as $\Im[z]$ diverges to infinity.
@@
Fix $\tau\in\mathbb{R}_{\geq 0}$ and assume that $M^{(\tau)}\in\mathcal{M}$ such that, for all $z\in\mathbb{H}$, $M^{(\tau)}(z)$ solves the RDEL (5). Then, $\|M^{(\tau)}(z)-M^{(\tau)}_{\infty}\|\to 0$ as $\Im[z]\to\infty$.
@@
We proceed block-wise. By Lemma 4.2, $\|M^{(\tau)}_{1,1}(z)\|\leq(\Im[z])^{-1}$. Consequently, $\|M^{(\tau)}_{1,1}(z)\|\to 0$ as $\Im[z]\to\infty$. Furthermore, it follows from Lemma A.6 and the assumed properties of the superoperator that $\Im[(\mathcal{T}^{(\tau)}_{1,1}(M^{(\tau)}))^{-1}]\succeq\Im[z](\mathcal{T}^{(\tau)}_{1,1}(M^{(\tau)}))^{-1}(\mathcal{T}^{(\tau)}_{1,1}(M^{(\tau)}))^{-\ast}$. Here, we use the notation $\mathcal{T}^{(\tau)}(M)=\mathbb{E}L-\mathcal{S}(M)-z\Lambda-i\tau I_{\ell}$ introduced in Equation 14. Taking the norm of both sides and rearranging, we obtain $\|(\mathcal{T}^{(\tau)}_{1,1}(M^{(\tau)}))^{-1}\|\leq(\Im[z])^{-1}$. Hence, $\|(\mathcal{T}^{(\tau)}_{1,1}(M^{(\tau)}))^{-1}\|\to 0$ as $\Im[z]\to\infty$. Furthermore, using the flatness of the superoperator, we have $\|\mathcal{S}_{1,2}(M^{(\tau)})\|\vee\|\mathcal{S}_{2,1}(M^{(\tau)})\|\vee\|\mathcal{S}_{2,2}(M^{(\tau)})\|\to 0$ as $\Im[z]\to\infty$. This implies that $\mathcal{T}_{1,2}(M^{(\tau)})\to\mathbb{E}B^{\top}$, $\mathcal{T}_{2,1}(M^{(\tau)})\to\mathbb{E}B$ and $\mathcal{T}_{2,2}(M^{(\tau)})\to\mathbb{E}Q-i\tau I_{d}$ as $\Im[z]\to\infty$. Since $\mathbb{E}Q-i\tau I_{d}$ is non-singular and the taking a matrix inverse is a continuous operation, $(\mathcal{T}_{2,2}(M^{(\tau)}))^{-1}\to M^{(\tau)}_{\ast}$ as $\Im[z]\to\infty$. Finally, using Equation 14, we conclude that $M^{(\tau)}(z)\to M^{(\tau)}_{\infty}$ as $\Im[z]\to\infty$. ∎
@@
Similarly to Lemma 4.4, it follows from lemmas A.9 and A.5 that $\|(L-z\Lambda-i\tau I_{\ell})^{-1}-\mathrm{diag}\{0_{n\times n},(Q-i\tau I_{d})^{-1}\}\|$ as $\Im[z]\to\infty$.
@@
Fix $\tau\in\mathbb{R}_{\geq 0}$ and assume that $M^{(\tau)}\in\mathcal{M}$ such that, for all $z\in\mathbb{H}$, $M^{(\tau)}(z)$ solves the RDEL (5). Then, there exists some constant $c\in\mathbb{R}_{>0}$ such that $\|M^{(\tau)}(z)-M^{(\tau)}_{\infty}\|\leq c(|z|-\kappa)^{-1}$ for all $z\in\{z\in\mathbb{H}\,:\,|z|>\kappa+c\kappa^{-1}\}$ with $\kappa:=2\|(\mathbb{E}Q-i\tau I_{d})^{-1}\|(\|\mathbb{E}B\|+(2\|(\mathbb{E}Q-i\tau I_{d})^{-1}\|)^{-1})^{2}+\|\mathbb{E}A\|+(2\|(\mathbb{E}Q-i\tau I_{d})^{-1}\|)^{-1}+s\|(\mathbb{E}Q-i\tau I_{d})^{-1}\|$.
@@
Fix $z\in\mathbb{H}$ with $|z|>\kappa$ and let $M^{(\tau)}\equiv M^{(\tau)}(z)$. For notational convenience, we denote $a=\|\mathbb{E}A\|$, $b=\|\mathbb{E}B\|$ and $m_{\star}=\|M_{\star}^{(\tau)}\|$. We will show that there exists $c\in\mathbb{R}_{>0}$ such that $\|M^{(\tau)}(z)-M_{\infty}^{(\tau)}\|\notin(c(|z|-\kappa)^{-1},\kappa]$ for every $z\in\mathbb{H}$ with $|z|>\kappa+c\kappa^{-1}$. By Lemma 4.4, $\|M^{(\tau)}(z)-M_{\infty}^{(\tau)}\|$ is in a neighborhood of the origin for every $z\in\mathbb{H}$ with $\Im[z]$ large enough. Since $z\mapsto\|M^{(\tau)}(z)-M_{\infty}^{(\tau)}\|$ is a continuous function on $\{z\in\mathbb{H}\,:\,|z|>\kappa+c\kappa^{-1}\}$, this will imply that $\|M^{(\tau)}(z)-M_{\infty}^{(\tau)}\|\in[0,c(|z|-\kappa)^{-1}]$ for every $z\in\{z\in\mathbb{H}\,:\,|z|>\kappa+c\kappa^{-1}\}$.
Suppose that $\|M^{(\tau)}-M^{(\tau)}_{\infty}\|\leq(2sm_{\star})^{-1}$ such that $\|M^{(\tau)}\|\leq\|M^{(\tau)}-M^{(\tau)}_{\infty}\|+\|M^{(\tau)}_{\infty}\|\leq(2sm_{\star})^{-1}+m_{\star}$. We consider the blocks separately using (14). By definition of $\mathcal{S}_{2,2}$, $\|\mathcal{S}_{2,2}(M^{(\tau)})\|\leq s\|M^{(\tau)}_{1,1}\|\leq(2m_{\star})^{-1}$. It follows from Lemma A.2 that
$\|(\mathcal{T}_{2,2}(M^{(\tau)}))^{-1}\|=\|(\mathbb{E}Q-i\tau I_{d})^{-1}\left(\mathcal{S}_{2,2}(M^{(\tau)})(\mathbb{E}Q-i\tau I_{d})^{-1}-I_{d}\right)^{-1}\|\leq 2m_{\star}.$
@@
$\|\mathbb{E}A-\mathcal{S}_{1,1}(M^{(\tau)})-\mathcal{T}_{1,2}(M^{(\tau)})(\mathcal{T}_{2,2}(M^{(\tau)}))^{-1}\mathcal{T}_{2,1}(M^{(\tau)})\|\leq a+(2m_{\star})^{-1}\\ +2m_{\star}(b+(2m_{\star})^{-1})^{2}.$
For $z\in\mathbb{H}$ with $|z|>\kappa\geq a+(2m_{\star})^{-1}+2m_{\star}(b+(2m_{\star})^{-1})^{2}$, it follows from Lemmas A.2 and 14a that $\|M^{(\tau)}_{1,1}\|\leq(|z|-\kappa)^{-1}$. We now turn our attention to $M^{(\tau)}_{2,2}$. Using Lemmas A.1 and 14b,
$M^{(\tau)}_{2,2}-M_{\star}^{(\tau)}=M^{(\tau)}_{2,2}\left(\mathcal{S}_{2,2}(M^{(\tau)})+\mathcal{T}_{2,1}(M^{(\tau)})(\mathcal{T}_{1,1}(M^{(\tau)}))^{-1}\mathcal{T}_{1,2}(M^{(\tau)})\right)^{-1}M^{(\tau)}_{\ast}.$
By Lemma A.2, $\|(\mathcal{T}_{1,1}(M^{(\tau)}))^{-1}\|\leq(|z|-a-(2m_{\star})^{-1}-sm_{\star})^{-1}$. Hence,
$\|M^{(\tau)}_{2,2}-M_{\star}^{(\tau)}\|\leq sm_{\star}((2sm_{\star})^{-1}+m_{\star})\|M^{(\tau)}_{1,1}\|\\ +m_{\star}((2sm_{\star})^{-1}+m_{\star})(b+(2m_{\star})^{-1})^{2}(|z|-\kappa)^{-1}.$
Plugging the bound for $\|M^{(\tau)}_{1,1}\|$ derived above and simplifying,
$\|M^{(\tau)}_{2,2}-M_{\star}^{(\tau)}\|\leq m_{\star}(s+(b+(2m_{\star})^{-1})^{2})((2sm_{\star})^{-1}+m_{\star})(|z|-\kappa)^{-1}.$
It only remains to treat $\|M^{(\tau)}_{1,2}\|$ and $\|M^{(\tau)}_{2,1}\|$, which we directly bound by
$\max\{\|M^{(\tau)}_{1,2}\|,\|M^{(\tau)}_{2,1}\|\}\leq 2m_{\star}(b+(2m_{\star})^{-1})\|M^{(\tau)}_{1,1}\|\leq 2m_{\star}(b+(2m_{\star})^{-1})(|z|-\kappa)^{-1}$
using (14c) and (14d). To summarize, we showed that for every $z\in\mathbb{H}$ with $|z|>a+(2m_{\star})^{-1}+2m_{\star}(b+(2m_{\star})^{-1})^{2}+sm_{\star}$, $\|M^{(\tau)}(z)-M^{(\tau)}_{\infty}\|\leq(2sm_{\star})^{-1}$ implies that
$\|M^{(\tau)}(z)-M^{(\tau)}_{\infty}\|\leq\|M^{(\tau)}_{1,1}(z)\|+\|M^{(\tau)}_{1,2}\|+\|M^{(\tau)}_{2,1}\|+\|M^{(\tau)}_{2,2}-M^{(\tau)}_{\star}\|\leq c(|z|-\kappa)^{-1}$